# Two-Stage Nuisance Function Estimation for Causal Mediation Analysis

### v2

@@ Appendix A Proofs
$T_{3}$ $=\frac{1}{\sqrt{L}}\sum_{\ell=1}^{L}\sqrt{m}\Big{\{}\mathbb{E}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\psi_{0}\Big{\}}$
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\left\{\mathbb{E}\left[\hat{\psi}^{mr}_{\ell}\right]-\psi_{0}\right\}$
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\Bigg{\{}\mathbb{E}\left[\left\{(1-A)\hat{\pi}_{1,\ell}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\}\right]$
$~{}~{}~{}~{}~{}~{}+\mathbb{E}\left[\left\{A\hat{\pi}_{2,\ell}(M,X)-(1-A)\hat{\pi}_{1,\ell}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\}\right]\Bigg{\}}$
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\Bigg{\{}\mathbb{E}\left[\left\{\mathbb{E}\left[1-A\mid X\right]\hat{\pi}_{1,\ell}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\}\right]$
$~{}~{}~{}~{}~{}~{}+\mathbb{E}\left[\left\{\mathbb{E}\left[A\mid M,X\right]\hat{\pi}_{2,\ell}(M,X)-\mathbb{E}\left[1-A\mid M,X\right]\hat{\pi}_{1,\ell}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\}\right]\Bigg{\}}$
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\Bigg{\{}\mathbb{E}\left[\left\{p\left(A=0\mid X\right)\hat{\pi}_{1,\ell}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\}\right]$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\mathbb{E}[\{p\left(A=1\mid M,X\right)\hat{\pi}_{2,\ell}(M,X)$

### v3

@@ Appendix A Proofs
$T_{3}$ $=\frac{1}{\sqrt{L}}\sum_{\ell=1}^{L}\sqrt{m}\Big\{E\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right]-\psi_{0}\Big\}$
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\left\{E\left[\hat{\psi}^{\text{2S}}_{\ell}\right]-\psi_{0}\right\}$
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\Bigg\{E\left[\left\{(1-A)\hat{\pi}_{1,\ell}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\}\right]$
$~~~~~~+E\left[\left\{A\hat{\pi}_{2,\ell}(M,X)-(1-A)\hat{\pi}_{1,\ell}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\}\right]\Bigg\}$
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\Bigg\{E\left[\left\{E\left[1-A\mid X\right]\hat{\pi}_{1,\ell}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\}\right]$
$~~~~~~+E\left[\left\{E\left[A\mid M,X\right]\hat{\pi}_{2,\ell}(M,X)-E\left[1-A\mid M,X\right]\hat{\pi}_{1,\ell}(X)\right\}\{\mu_{1}(M,X)-\hat{\mu}_{1,\ell}(M,X)\}\right]\Bigg\}$
$=\frac{\sqrt{m}}{\sqrt{L}}\sum_{\ell=1}^{L}\Bigg\{E\left[\left\{p\left(A=0\mid X\right)\hat{\pi}_{1,\ell}(X)-1\right\}\{\mu_{2}(X)-\hat{\mu}_{2,\ell}(X)\}\right]$
$~~~~~~~~~~~~~~~+E[\{p\left(A=1\mid M,X\right)\hat{\pi}_{2,\ell}(M,X)$